OpenStudy (anonymous):

Consider this system of linear equations: y − x = -13 -4x + 3y = -51

OpenStudy (anonymous):

choices (-12,1) (-12,-1) (12,-1)

OpenStudy (anonymous):

Hi! I think I can help you with this :)

OpenStudy (anonymous):

First of all, any ideas about how to get started, or are you completely stuck?

OpenStudy (anonymous):

im completelty stuck on the lesson its self :(

OpenStudy (anonymous):

Okay, no problem!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I can think of two possible ways to answer this question. One way would be to try and solve the equations yourself. Alternatively, we could just try each of the multiple choice answers and see which one works.

OpenStudy (jameshorton):

mines say 0,0

OpenStudy (anonymous):

Any preference?

OpenStudy (anonymous):

im thinking that it is -12,1

pooja195 (pooja195):

Id start with adding x to both sides y − x +x= -13+x y=-13+x The subtitute the y value in place of 3y -4x + 3(-13+x) = -51 <---solve for x

OpenStudy (anonymous):

that looks confusing

pooja195 (pooja195):

Not at all :) \[\huge~\rm~-4x + 3(-13+x) = -51 \] Distribute the 3 \[\huge~\rm~ 3(-13+x) \]

OpenStudy (anonymous):

-39+3x

pooja195 (pooja195):

Good -39+3x \[\huge~\rm~-4x -39+3x = -51 \] Combine like terms

OpenStudy (anonymous):

im confused

pooja195 (pooja195):

See the x's ? Combine them since they are like terms -4+3=?

OpenStudy (anonymous):

-1x

pooja195 (pooja195):

Good now add −x−39+39=−51+39 -51+39=?

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

right?

pooja195 (pooja195):

its just 12

OpenStudy (anonymous):

no bevause 51 is negative.

pooja195 (pooja195):

a negative divided by a negative is a postivie

OpenStudy (anonymous):

but its not divided its added

pooja195 (pooja195):

Answer what i asked you -51+39=?

OpenStudy (anonymous):

-12